hi, my son asked me some questions on algebra but i myself was confused. hope u can help me to help him with these.write these follwing expressions in the simplest form:
1. b+7+2b+10
2. (3d+5)+(d-2)
3. 3m-(-m)

Answer by Earlsdon(6294)

Simplify:
1. b+7+2b+10 Add the like-terms together. The like-terms are: b & 2b and 7 & 10. So we get:
(b+2b) + (7+10) after adding, we have:
3b+17
2. (3d+5)-(2d-5) Again, add together the like-terms which are: 3d & -2d and 5 & +5 Notice the change in sign (From - to +) on the last 5 after you remove the parentheses. 3d+5-2d+5
(3d-2d) + (5+5) = d+10
3. 3m-(-m) Remember that to subtract a number you add its opposite (same number but opposite sign).
3m+m = 4m